# Infinigen-Sim: Procedural Generation of Articulated Simulation Assets

### v2

@@ 5.2 Reinforcement Learning Generalization
    Push Down Toaster Lever. The robot is tasked with pushing down the lever of the toaster. Only toasters with a single lever are considered for this task. We evaluate on 5 unseen PartNet toasters and compare RL policies where one is trained on the remaining 8 PartNet toasters while the other is trained on 8 PartNet and 14 Infinigen-Sim toasters.
@@
    Push Door with Handle. The robot needs to push the door open by first rotating the door handle clockwise by more than 40 degrees. Only single doors with a lever handle are considered for this task. We evaluate on 3 PartNet doors with handles and compare RL policies where one is trained on the remaining 3 PartNet doors while the other is trained on 3 PartNet and 44 Infinigen-Sim doors.
@@
    Open Fridge Door. The robot is tasked with grasping the handle and opening the fridge door. Only refrigerators with a single door are considered for this task. We evaluate on 5 PartNet fridges and compare RL policies where one is trained on the remaining 5 PartNet fridges while the other is trained on 5 PartNet and 44 Infinigen-Sim fridges.
@@
We train PPO [46] policies using GPU-parallel heterogeneous simulation across 64 environments in ManiSkill3 [2]. Simulator environments using Infinigen-Sim assets ran at speeds indistinguishable from those using PartNet-Mobility. For all tasks, the policy takes as input an RGB image from a third-person view and its end-effector pose with no additional privileged information. The output at each time step is a delta pose command for operational space control (OSC). Refer to the appendix for details.
Figure 4 shows the learning curve (averaged over 4 runs) during evaluation on unseen PartNet assets. It is clear from this figure that training policies with a combination of Infinigen-Sim and PartNet assets results in better generalization than training with only PartNet assets. PartNet-Mobility assets provide limited visual and geometric diversity across these categories. Additionally, misaligned joints (e.g. an incorrect pivot position for a door handle) may make tasks harder to complete and hinder policy learning. In contrast, our tool provides dynamically accurate assets with both great visual and geometric diversity. Our assets are also able to cover a wider distribution of values for important parameters (see Figure 3) that impact a robot’s trajectory. Through these RL experiments, we claim that our tool complements existing datasets and can boost generalization capabilities to unseen model instances.
Figure 5: An imitation learning policy trained using Infinigen-Sim assets (left) achieves a 7/10 success-rate on opening a real door (right).
5.3 Sim-to-Real Transfer
@@
We show that Infinigen-Sim assets are effective in sim-to-real transfer. We created a real-world version of the Push Door with Handle task using a Franka Panda robot (Fig. 5 right). For simplicity, we train imitation learning models using expert simulation trajectories. We set up an IsaacGym [47] environment for various door assets (Fig. 5 left). As we have the ground truth joint axis and origin, we rollout scripted trajectories and filter successful ones as the demonstration dataset. The robot takes an RGB image from its on-the-shoulder camera and the end effector pose as inputs. We train ACT [48] on 1k+ trajectories with randomized lighting, joint initialization, and camera views. For comparison, we curate two datasets including one with 300+ Infinigen-Sim assets and another with all 6 PartNet assets.
We evaluate the zero-shot performance on a real world door. For each trial, we randomly initialize the robot’s pose. We observe that policies trained with Infinigen-Sim assets achieve 7 out of 10 successes in the real world, while those trained with PartNet assets fail on all 10 trials. This increase in performance is likely a result of the diversity and quantity of assets generated with our tool. These experiments show that our tool has the ability to create articulated assets diverse enough to facilitate sim-to-real transfer.
6 Conclusion
@@
We introduce Infinigen-Sim, a tool for creating procedural generators for articulated simulation-ready assets. These assets are kinematically diverse, photorealistic, and have accurate joint configurations. Our tool also gives users fine-grained control over their assets. We use this system to create expressive generators for 5 common articulated object categories. Our experiments demonstrate that objects from these generators can improve model generalization on both perception and robot learning tasks in simulation and the real world.
7 Limitations
Our work currently does not consider material physical properties such as friction or restitution. With regards to our custom joint nodes, there exist minor restrictions in the way an object’s node graph should be constructed. However, we note these restrictions do not limit the types of articulated assets that can be produced, merely the structure of the graph. We hope to continue developing our custom joint node and lift these restrictions to make our tool even more user-friendly.
For movable part segmentation, we exclusively evaluate our models on data generated using PartNet assets alone. This means certain part categories may be limited in diversity and quantity. Future work may explore the benefit of using Infinigen-Sim assets when evaluating on annotated real world images of articulated objects.
Acknowledgments
@@
A.1 Example Procedural Assets
In this section, we present several examples of assets created using our custom generators. The distributions shown reflect the default settings but can be customized as needed.
Figure 6: Procedurally generated articulated doors with handles.
Figure 7: Procedurally generated articulated toasters with levers, buttons, and knobs.
Figure 8: Procedurally generated articulated refrigerators with internals/external drawers and shelves.
Figure 9: Procedurally generated articulated dishwashers with handles, racks, buttons, and knobs.
Figure 10: Procedurally generated articulated lamps with many different types of switches.
A.2 Example Articulations
In this section, we showcase minimal examples of how our custom joint nodes can be used to create different types of articulations within Blender. These articulated assets directly translate to simulation using our native exporters for the URDF, USD, and MJCF file types.
Figure 11: Simple Revolute Joint: Demonstrates a rod rotating about a pivot point. Our nodes include togglable debugging features, such as axis visualization, to assist users in verifying joints.
Figure 12: Simple Prismatic Joint: Demonstrates an articulated button using our custom joint node.
Figure 13: Duplicating Jointed Bodies: Demonstrates duplication of multiple articulated knobs. Instead of manually defining a joint for each knob, this node automatically replicates the articulated body at a set of points (a grid in this example).
Figure 14: Connecting Multiple Jointed Bodies: Demonstrates connecting multiple bodies together. We first joint the two rods together, followed by jointing the combined body with the base.
Figure 15: Jointing to the Same Body: Demonstrates jointing two rods to the same sphere. First, one rod is jointed to the sphere to form a body. Then, the second rod is jointed to this body by attaching it to the top-most parent object (i.e. the sphere).
Figure 16: Multi-Jointed Bodies: Demonstrates a multi-jointed articulated cap. Combining a hinge and slide joint can achieve a screw joint commonly used for for assets such as water bottles.
A.3 List of Interpretable Degrees of Freedom
Articulated Object DOF Interpretable Parameters
Toaster 14 Dimension of Lever handle, Slot Width, Slot Length, Slot Depth, Toaster Length, Knob Vertical Location, Knob Horizontal Location, Knob Size, Circular Button Size, Square Button Width, Inter-Button Distances, Button Horizontal Offset, Button Vertical Offset, Protrusion parameter
Fridge 32 Size, Wall Thickness, Body Outer Roundness, Body Inner Roundness, Door Handle Margin, Door Shelf Size, Door Shelf Thickness, Door Shelf Num, Door Shelf Margin, Door Handle Top Size, Door Handle Top Thickness, Door Handle Top Roundness, Door Handle Support Size, Door Handle Support Margin, Door Left Margin, Door Right Margin, Door Upper Margin, Door Lower Margin, Shelf Margin, Shelf Thickness, Shelf Board Margin, Drawer Height, Drawer Wall Thickness, Drawer Handle Margin, Drawer Handle Top Size, Drawer Handle Top Thickness, Drawer Handle Top Roundness, Drawer Handle Support Size, Drawer Handle Support Margin, Drawer Body Roundness, Drawer Slide Roundness, Drawer Inner Roundness
Dishwasher 13 Depth, Width, Height, Door Thickness, Rack Radius, Rack Height, Rack Depth, Handle Radius, Handle Position, Number of Racks, Density of Supports in Rack, Button Position, Handle Curvature
Lamp 29 Pull String Radius, Pull String Base Height, Pull String Length, Button Base Size, Button Size, Button Height, Twist Button Base Size, Twist Button Size, Twist Button Height, Twist Button Twister height, Switch Base Size, Switch Size, Switch Curvature, Button X-Location, Length of Bar 1, Location of Bar 2 joint on Bar 1, Length of Bar 2, Location of Bar 3 joint on Bar 2, Radius, Height, Radius of Base, Base Height, Number of Sides on Base, Shade Height, Rack Height, Top Radius, Bottom Radius, Rack Thickness, Number of Sides on Shade
Door 39 Width, Height, Depth, Panel margin, Bevel Width, Shrink Width, Door Frame Width, Handle Height, Push Bar Length, Push Bar Thickness, Push Bar Aspect Ratio, Push Bar Height Ratio, Push Bar Length Ratio, Push Bar End Length Ratio, Push Bar End Height Ratio, Push Bar Overall Z-Offset, Knob Radius, Knob Base Radius, Knob Middle Radius, Knob Central Radius, Knob Depth, Knob Middle Depth, Lever Radius, Lever Middle Radius, Lever Depth, Lever Middle Depth, Lever Length, Lever Type, Pull Handle Size, Pull Handle Depth, Pull Handle Width, Pull Handle Extension, Pull Handle Bevel Width, Pull Handle Pull Radius, Pull Handle Bevel Side Length, Louver Width, Louver Margin, Louver Size, Louver Angle
Table 3: Toaster, refrigerator, dishwasher, lamp, and door generators with interpretable degrees of freedom and named parameters
A.4 Experiment Setup of Movable Part Segmentation
@@
In order to run the experiments as described in Sec. 5.1, we first generate images using and ground truth labels and segmentations for both the PartNet-Mobility and Infinigen-Sim assets. Each image is of size $1024\times 1024$ to ensure that smaller movable parts are visible. Similar to Xiang et al. [18], images are taken from a camera positioned randomly on the upper hemisphere of the asset with the scene having ambient lighting and fixed directional lighting. We also randomly sample the field of view. Finally, we randomize the joint values so that the same asset appears in different poses across images. For assets such as dishwashers where interior racks may intersect with exterior bodies, we sample joint values within a range to try to avoid this. All images are rendered using ray-tracing with Maniskill3 [2].
@@
In order to curate the evaluation dataset for the movable part segmentation experiments, we first randomly sample PartNet-Mobility assets such that there exists at least one asset for each considered movable part. To ensure fairness and valid evaluation, we then remove invalid assets from the evaluation set based on the following criteria:
  1. 1.
    Asset includes parts that should be movable but are not (i.e. the mesh of a button is present, but the button is not articulated).
  2. 2.
@@
  3. 3.
    Asset contains incorrect face normals, resulting in transparent parts.
We then subsample the evaluation dataset such that we have 25% of assets from each model category, ensuring at least one asset remains for each movable part. Both invalid and subsampled assets are moved back to the PartNet-Mobility training dataset.
During training, we use the following hyperparameters to train our Mask R-CNN model:
@@
Learning Rate 1e-4
Weight Decay 1e-5
Evaluation Mask Treshold 0.5
@@
Epochs 10
Table 4: Mask R-CNN Training Hyperparameters
A.5 Overall Results of Movable Part Segmentation
In this section, we provide the overall results for the movable part segmentation experiments. We observe an overall increase in performance when training our Mask R-CNN model on a combination of both PartNet-Mobility and Infinigen-Sim assets. We attribute this boost in performance to the increase diversity introduced when training on Infinigen-Sim assets. This diversity likely allows the model to be more robust to unseen objects.
Dataset $mAP$ $mAP_{50}$ $mAP_{75}$ $mAP_{small}$ $mAP_{medium}$ $mAP_{large}$
Partx1 48.23 63.22 50.14 9.09 30.03 62.46
Partx2 48.46 63.42 50.29 9.37 29.57 62.93
Partx1+Inf 50.13 64.92 51.97 9.92 32.10 62.27
Table 5: Movable part segmentation mAP scores for all experiments averaged across three seeds. We train each model for 50 epochs.
A.6 Experiment Setup of RL Generalization
We train RL agents on Push Down Toaster Lever, Push Door with Handle, and Open Fridge Door to show that Infinigen-Sim assets create a better asset distribution for generalizable robot policies on new instances. In all our experiments, we only use assets from PartNet-Mobility [18] for evaluation. For fair comparison, all policies are trained with 64 GPU-based parallel environment with instances sampled from the training distributions, i.e., PartNet only v.s. PartNet + Infinigen-Sim. In Figure 4, we plot the learning curve of evaluation success rate averaged over 4 seeds.
Push Down Toaster Lever. The robot is tasked with pushing down the lever of the toaster. Only toasters with a single lever are considered for this task. We evaluate on 5 unseen PartNet toasters and compare RL policies where one is trained on the remaining 8 PartNet toasters while the other is trained on 8 PartNet and 14 Infinigen-Sim toasters.
In RL training, we use the reward to encourage the fingertip to approach the lever, to push down the lever, and to penalize on the action scale. The task is considered as a success if the lever is pushed more than 50% of its joint range. The policy takes an on-the-shoulder, third-person RGB image (Figure 4) + fingertip pose as inputs and outputs a 6-dim delta pose change. The action commands the Operation Space Controller (OSC).
Push Door with Handle. The robot needs to push the door open by first rotating the door handle clockwise by more than 40 degrees. Only single doors with a lever handle are considered for this task. We evaluate on 3 PartNet doors with handles and compare RL policies where one is trained on the remaining 3 PartNet doors while the other is trained on 3 PartNet and 44 Infinigen-Sim doors.
In RL training, we use the reward to encourage the fingertip to approach the handle, to rotate the handle, and to push the door to open. We consider a trial as a success if the door is opened by at least 15 degrees. The policy takes an on-the-shoulder, third-person RGB image (Figure 4) + fingertip pose as inputs and outputs a 6-dim delta pose change. The action commands the Operation Space Controller (OSC).
Open Fridge Door. The robot is tasked with grasping the handle and opening the fridge door. Only refrigerators with a single door are considered for this task. We evaluate on 5 PartNet fridges and compare RL policies where one is trained on the remaining 5 PartNet fridges while the other is trained on 5 PartNet and 44 Infinigen-Sim fridges.
In RL training, we use the reward to encourage the fingertip to approach the door and to open the door. We consider the trial as a success if the door is opened by at least 15 degrees. The policy takes an on-the-shoulder, third-person RGB image (Figure 4) + fingertip pose as inputs and outputs a 6-dim delta pose change. The action commands the Operation Space Controller (OSC). Noticing that, we do not require to use the handle to successfully open the door.
In our results (Fig. 4), we show that policies trained with both Infinigen-Sim and PartNet assets can have a better generalization performance and more stable learning curve than the policy trained with PartNet assets only. However, the current success rate is still low to solve the tasks. We leave this as future work to train robust and better policies for these tasks with potentially more Infinigen-Sim assets.
A.7 Experiment Setup of Sim-to-Real Transfer
We apply a procedural approach to generate expert trajectories. Giving handle assets, we pre-sample candidate pre-grasps and grasp poses according to the handle position and its bounding box. Moreover, we use CuRobo to solve IK of pre-grasp poses as initialization of trajectories and use OSC to generate a cartesian linear trajectory to the grasp pose. After finger closure, we generate a circular cartesian motion to rotate the handle along the handle joint axis. The axis direction and its origin is known given the asset annotations. After the rotation is more than 45 degrees, we generate door opening circular motion, given the door joint position and origin. As this procedural pipeline may not work for some grasp poses and door assets, we filter the trajectories with success criteria and use the curated trajectories as the imitation learning dataset. All our trajectories are generated and rendered in IsaacGym [47], with its default rendering engine.
@@ A.7 Experiment Setup of Sim-to-Real Transfer
The imitation learning policy outputs target 6-dof poses in the robot base and 1-dim for gripper action, instead of delta motion [48]. Here, the input only consists the on-the-shoulder, third-person view RGB image and its fingertip pose (Fig. 5). The training configuration follows [48]. On a real Franka robot, we use DEOXYS to command the real robot with an OSC controller for the absolute target pose. We set a relative stiff PD value as the task requires certain strength to solve.

### v3

@@ V-B RL Generalization
    Push Door with Handle. The robot is required to rotate the handle clockwise by 0.3 radians and push the door open. We limit the scope to single, push doors with a lever handle on the left.
@@
    Push Down Toaster Lever. The robot is tasked with pushing down the lever of the toaster. Only toasters with a single lever are considered for this task.
@@
    Open Fridge Door. The robot is expected to grasp the handle and open the fridge door by at least 9 degrees ($\pi/20$ radians). Only refrigerators with a single door with a handle on the left are considered for this task.
@@
To ensure fairness, our evaluation sets consists of a mixture of 5 unseen Infinigen-Articulated assets and 5 unseen PartNet-Mobility assets. We train PPO [46] policies in ManiSkill3 [1] using GPU parallelized heterogeneous simulation across 128 environments, where each environment may contain a different asset. The PPO policy takes as input its end-effector pose and an RGB image rendered by the default ManiSkill3 renderer from a third-person view with no additional privileged information. The output at each time step is a delta pose command for operational space control (OSC). We refer the reader to the appendix for details.
Figure 5: Left: Environment setups for each task. Right: Average success-once performance over 5 seeds with bounds showing $\pm 0.5$ standard deviation.
Our results demonstrate a 2.86-fold average improvement in success rate across all three tasks (during the last 1 million timesteps) when using exclusively Infinigen-Articulated assets over PartNet-Mobility. We attribute this performance increase to the scale, quality, and diversity of Infinigen-Articulated assets. We hypothesize that the limited number of PartNet-Mobility assets, together with potentially misaligned joints and poor geometries, significantly increases the difficulty of training effective policies. In contrast, our tool provides dynamically accurate assets with both great visual and geometric diversity. However, it is worth noting that during the earlier stages of training, for the door and fridge opening tasks, the PartNet-Mobility policy performed on par or better than policies trained using Infinigen-Articulated. This suggests that realizing the benefits of more diverse assets may require longer training times.
V-C Sim-to-Real Transfer
@@
We also show that Infinigen-Articulated assets are effective in sim-to-real transfer. We create real-world versions of the door opening, toaster pushing, and fridge opening tasks using a Franka Panda robot (see Fig. 6). For each task, we collect expert simulation trajectories using cuRobo [47] in an IsaacGym [48] environment. We randomize environment lighting, joint initialization, object position, and camera views. As we have the ground truth joint axis and origin, we roll out scripted trajectories and filter successful ones as the demonstration dataset. The robot takes an RGB image from its on-the-shoulder camera and the end effector pose as inputs. We then train an ACT [49] policy on this dataset and transfer it zero-shot to a physical robot.
Figure 6: Real world environment setup. From left to right: push open door, push toaster lever, and pull open fridge door.
Open Door Push Toaster Pull Fridge
PartNet-Mobility [17] 3/30 0/30 0/30
Infinigen-Articulated (Ours) 18/30 11/30 10/30
TABLE II: Zero-shot sim-to-real results.
We observe that policies trained using Infinigen-Articulated perform significantly better across all three tasks (see Table II). Qualitatively, we notice that policies trained using our assets tend to be smoother and fail less aggressively. For instance, failures in the fridge task by the Infinigen-Articulated policy typically involved either moving toward the handle and nearly grasping it or opening the door less than the required 45 degree threshold. In contrast, policies trained on PartNet-Mobility often missed the handle altogether. Similar to the RL experiments, we attribute this success to the increased diversity and quantity of assets generated with our tool. We postulate that it is this diversity that helps bridge the sim-to-real gap.
VI Limitations and Future Work
Joints are currently modeled independently. While coupling them (e.g., pressing the pedal of a trash can to open the lid) is a relatively straightforward extension, we leave this for future work. Additionally, Infinigen-Articulated currently uses CoACD [43] to generate collision meshes for individual parts. Depending on the level of detail of the convex decomposition, this can potentially slow simulation. That being said, a key advantage of Infinigen-Articulated is that we know how each part is constructed via its node graph. We leave it as future work to determine how to generate a corresponding set of primitives to serve as the asset’s collision mesh using the node graph, potentially improving simulation speed.
VII Conclusion
@@
We introduce Infinigen-Articulated, a toolkit to generate procedural, articulated, simulation-ready assets. We use the system to provide procedural generators for 18 articulated object categories. These assets are kinematically diverse, photorealistic, follow consistent naming conventions, and have accurate joint configurations. Infinigen-Articulated provides users fine-grained control over their assets including sampling from physical parameters to bridge the sim-to-real gap. Our experiments demonstrate that objects from these generators can improve model generalization on both perception and robot learning tasks in simulation and the real world.
ACKNOWLEDGMENT
@@
VII-A Articulated Parts
Table III provides a list of articulated parts for each of our procedural generators. Certain parts may have their own procedural parameters including style, count, positioning, dimensions, etc.
Object Category Articulated Parts
Door Door, handle
Cooktop Burner dial
Lamp Lamp arms, shade, switch/button/dial/pull chain
Box Box flaps
Soap dispenser Nozzle
Oven Door, racks, drawer
Toaster Lever, buttons, knobs
Cabinet Doors
Drawer Doors
Pepper grinder Top bulb/handle
Refrigerator Doors, internal drawers, external drawers
Trashcan Lid, pedal
Dishwasher Door, buttons, racks, knobs
Faucet Spout, handles
Microwave Door, internal plate, eject button, dials
Window Window panes, locks
Plier Plier hand
Handle Main handle, lock
TABLE III: Provided object categories and their articulations.
VII-B Moveable Part Segmentation Experiment Details
@@
To curate a dataset, we generate images using segmentation ground truth for both the PartNet-Mobility and Infinigen-Articulated assets. Each image is of size $1024\times 1024$ to ensure that smaller articulated parts are visible. Similar to [17], images are taken from a camera positioned randomly on the upper hemisphere of the asset with the scene having ambient lighting and fixed directional lighting. We also randomly sample the field of view. Finally, we randomize joint values such that the same asset appears in different poses across images. For assets such as dishwashers where interior racks may intersect with exterior bodies, we sample joint values within a range to try to avoid this situation. All images are rendered using ray-tracing with Maniskill3 [1].
@@
We then randomly sample PartNet-Mobility assets for our evaluation set such that there exists at least one asset for each considered movable part. To ensure fairness, we remove invalid assets from the evaluation set based on the following criteria:
  • Asset includes parts that should be movable but are not.
@@
  • Asset contains incorrect face normals.
We subsample the evaluation dataset such that we have 25% of assets from each model category, ensuring at least one asset remains for each movable part. Remaining assets are moved back to the PartNet-Mobility training dataset. Hyperparameters for training are shown in Table IV.
@@
Learning Rate $1\times 10^{-4}$
Weight Decay $1\times 10^{-5}$
Evaluation Mask Threshold 0.5
@@
TABLE IV: Mask R-CNN training hyperparameters.
VII-C RL Generalization Experiment Details
We define our reward functions as shown below where $p_{x}$ is the position of the $x$, $\hat{q}_{x}$ is the normalized joint configuration, $\lambda_{a}=0.1$ is the action penalty scale, and $T(x)=1-\tanh(x)$. If the task is successful, we assign an $R_{\text{success}}$ reward greater than the maximum reward defined by the functions below.
$r_{\text{door}}(s,a)$ $=T\!\left(10\,\lVert p_{\text{eef}}-p_{\text{handle}}\rVert\right)+4T\!\left(3(1-\hat{q}_{\text{handle}})\right)$
$\qquad+5T\!\left(2(1-\hat{q}_{\text{door}})\right)-\lambda_{a}\lVert a\rVert$
$r_{\text{toaster}}(s,a)$ $=T\!\left(10\,\lVert p_{\text{eef}}-p_{\text{lever}}\rVert\right)+2T\!\left(3(1-\hat{q}_{\text{lever}})\right)-\lambda_{a}\lVert a\rVert$
$r_{\text{fridge}}(s,a)$ $=T\!\left(5\,\lVert p_{\text{eef}}-p_{\text{handle}}\rVert\right)+2(1-(1-\hat{q}_{\text{door}})^{4})-\lambda_{a}\lVert a\rVert$
Hyperparameters for all RL experiments are shown in Table V.
Hyperparameter Value
Learning Rate $1\times 10^{-4}$
Gamma 0.95
Number of Parallel Environments 128
TABLE V: PPO training hyperparameters.
VII-D Sim-to-Real Transfer Experiment Details
We generate expert trajectories for each task using a Franka Panda robot in IsaacGym [48]. During each step, we store both the robot’s fingertip’s pose and an RGB image from an on-the-shoulder, third-person camera (rendered using IsaacGym’s default renderer). To generate trajectories, we sample candidate pre-grasps and grasp poses according to the handle position for the door and fridge tasks and the lever position for the toaster task. An operation space controller (OSC) is then used to generate a trajectory. We design specific paths for each task:
  • Pushing open doors: linear motion to the handle, circular motion to rotate the handle 45 degrees (based on handle–joint distance), and circular motion to push the door open (based on handle–hinge distance).
  • Pushing toaster lever: linear motion to the lever.
  • Pulling open fridge: linear motion to the handle, grasp, and circular motion to pull the door open (based on handle–hinge distance).
@@ VII-D Sim-to-Real Transfer Experiment Details
We train our policy on fingertip poses and RGB images. The policy outputs a target 6-DOF pose for the arm in the robot base frame and a single value for the gripper action instead of delta motion [49]. The training configuration follows [49]. We train multiple policies with different hyperparameters for 300 epochs and select the checkpoint with the lowest validation loss. The final hyperparameters are described in Table VI. We use Deoxys [50] to command the physical robot with an OSC controller for the absolute target pose. We set a relative stiff PD value.
Dataset Hyperparameter Door Toaster Fridge
Ours Learning Rate $1\times 10^{-4}$ $5\times 10^{-5}$ $1\times 10^{-4}$
KL Weight 1 10 100
Temp. Agg. $\checkmark$ $\checkmark$
PartNet Learning Rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $5\times 10^{-5}$
KL Weight 10 10 10
Temp. Agg. $\checkmark$
TABLE VI: Final ACT policy hyperparameters for each task on Infinigen-Articulated and PartNet-Mobility.
VII-E Comment on Robot Safety
For the real-world fridge opening task, the handle occasionally slipped into a gap between the robot’s fingers, causing it to get stuck. When the robot pulled back, this created excessive torque on the gripper and raised safety concerns. To address this, we add padding between the door and the base of the fridge as well as on the gripper to close the gap.